### v2

# From Intentions to Techniques: A Comprehensive Taxonomy and Challenges in Text Watermarking for Large Language Models

Abstract
With the rapid growth of Large Language Models (LLMs), safeguarding textual content against unauthorized use is crucial. Watermarking offers a vital solution, protecting both - LLM-generated and plain text sources. This paper presents a unified overview of different perspectives behind designing watermarking techniques through a comprehensive survey of the research literature. Our work has two key advantages: (1) We analyze research based on the specific intentions behind different watermarking techniques, evaluation datasets used, and watermarking addition and removal methods to construct a cohesive taxonomy. (2) We highlight the gaps and open challenges in text watermarking to promote research protecting text authorship. This extensive coverage and detailed analysis sets our work apart, outlining the evolving landscape of text watermarking in Language Models.
[…]
theoTheorem
From Intentions to Techniques: A Comprehensive Taxonomy and Challenges in Text Watermarking for Large Language Models
Harsh Nishant Lalai [Uncaptioned image]          Aashish Anantha Ramakrishnan [Uncaptioned image] Raj Sanjay Shah [Uncaptioned image]          Dongwon Lee [Uncaptioned image] Birla Institute of Technology and Science, Pilani [Uncaptioned image] The Pennsylvania State University [Uncaptioned image]          Georgia Institute of Technology [Uncaptioned image]
footnotetext: Email: f20212665@goa.bits-pilani.ac.in, {aza6352, dongwon}@psu.edu, rajsanjayshah@gatech.edu
1 Introduction
Large Language Models (LLMs) can mimic human-like comprehension and text generation Zheng et al. (2024). Consequently, judging whether a text is authored by a human or generated by an LLM is challenging. This issue is highlighted by the recent lawsuit of The New York Times against OpenAI and Microsoft concerning the use of their articles as training data for AI models, emphasizing the need for effective methods to identify and safeguard digital content ownership New York Times Company (2023).
Text Watermarking provides key solutions to protect intellectual property rights, identify ownership, and keep track of digital content. These techniques embed imperceptible signals or identifiers within digital text documents, which are then used to track the document’s origins Jalil and Mirza (2009); Kamaruddin et al. (2018). In particular, they aid in tracking the different production sources of text, both human-written and LLM-generated, helping prevent their unauthorized use without the owner’s consent.
Given this increasing research focus on watermarking techniques, it is important to review various methods, their applications, strengths and limitations. This includes systematically categorizing current research literature and highlighting key open challenges. The following contributions of our work distinguish it from previous surveys:
  • Taxonomy Construction: We seek to help future researchers navigate text-watermarking by categorizing various techniques and methods. Unlike traditional surveys, our paper aims to use the constructed taxonomy to provide an up-to-date list of research challenges for the text watermarking field instead of the most up-to-date survey of the field. For this task, we focus on application-driven intentions, evaluation data sources, and watermark addition methods. We also enlist potential adversarial attacks against these methods to caution readers.
  • Open Challenge Identification: Next, we describe open challenges and gaps in current research efforts. These span rigorous testing of methods against diverse de-watermarking attacks, the establishment of standardized benchmarks for appropriate method efficacy comparison, expansion of the downstream NLP tasks used for evaluation, understanding how watermarking impacts language model factuality and utility, the interpretability of watermarking techniques by detailed descriptions and visual aids, developing personalized watermarking techniques, and lastly, collective watermark creation in multi-agent settings.
Our work aims to enable researchers to recognize emerging trends and areas for improvement in text watermarking research. We facilitate this goal by creating a systematic and comprehensive taxonomy of text watermarking.
2 Taxonomy of Text Watermarking
To help researchers navigate the field of text watermarking, we cluster various techniques and methods based on key commonalities. For this categorization, we focus on intentions that are application-driven, data sources for model evaluation, watermark addition methods, and method-specific adversarial attacks. In our taxonomy creation, we allow techniques to belong to multiple categories and show how different techniques relate across multiple dimensions, making it easier to navigate the field.
2.1 Intention
Methods for embedding textual identifiers to watermark differ based on a user’s desired features, the user’s role (developer vs end-user, etc.), and primary application-driven needs. We categorize watermarking techniques based on the end user’s intention into three types: Text Quality, Output Distribution, and Model Ownership Verification.
2.1.1 Text Quality
Maintaining the quality and utility of the generated text post-watermarking is a desired goal of any watermarking methodology. However, research works differ on definitions of quality and mainly proxy output quality with (1) generation perplexity (uncertainty) and (2) semantic relatedness of watermarked and un-watermarked generations.
Minimizing impact on Perplexity
A model’s confidence in its generations, measured through the weighted sum of individual token log probabilities in a sequence is known as Perplexity. A lower perplexity indicates that the model is more confident and accurate in its predictions, while a higher perplexity suggests more significant uncertainty and less accurate predictions. Perplexity is the only intrinsic measure of model uncertainty Magnusson et al. (2023), and thus, a popular measure of quality among researchers.
Figure 1: An example of green-red list grouping of texts Kirchenbauer et al. (2023).
Watermarking techniques like using green-red list rules (refer to figure 1) trade-off the ability to detect LLM-generated text and the utility of the output text. For a given text, the greater the proportion of green tokens from the total tokens, the lesser the chance of the text being written by humans. A parallel aim is to reduce all of the other "generic" text’s perplexity while enforcing a more frequent generation of white-listed "green" words. Controlling entropy levels ensures that watermarked text maintains a quality similar to non-watermarked text. Studies have effectively operationalized this technique in diverse ways for authorship detection while maintaining high text quality Kirchenbauer et al. (2023); Zhao et al. (2023a); Takezawa et al. (2023). For example, soft watermarking promotes green list use for high-entropy (rare) tokens while minimally affecting low-entropy (common) tokens Kirchenbauer et al. (2023); Lee et al. (2023); Ren et al. (2024), ensuring that watermark is undetectable (soft) to an observer. In another example,Takezawa et al. (2023) recommend a lower watermark strength for longer texts for quality. Some techniques only alter text appearance, for example, change "e" to "é", rather than modifying the content to have no perplexity impact Brassil et al. (1995); Por et al. (2012); Sato et al. (2023).
Table 1: Overview of watermarking techniques using semantic relatedness. Struct: Maintains Structure, Word repl: Synonym/ Spelling based word replacement techniques, Dep. trees: Dependency trees, Syn. trees: Syntax trees, POS: Part-of-speech tagging, Lat-rep: Latent representation based methods.
Work Struct Word Dep. Syn. POS Lat.
repl. trees trees rep.
Topkara et al. (2006b)
Meral et al. (2009)
Abdelnabi and Fritz (2021)
Yang et al. (2022)
He et al. (2022a)
He et al. (2022b)
Yoo et al. (2023a)
Yang et al. (2023b)
Munyer and Zhong (2023)
Fu et al. (2024)
Hoang et al. (2024)
Semantic Relatedness
Refers to how closely words, phrases, or sentences of the watermarked output are similar to the original clean output. One way of maintaining input semantics is by embedding both input and output sentences into a semantic space and minimizing the distance between them Abdelnabi and Fritz (2021); Zhang et al. (2023). Yang et al. (2022) use the BERT model to suggest substitution candidates, while other works use synonyms and spelling replacements to have minimum impact on semantic relatedness. Fu et al. (2024) use the input context to extract semantically related tokens, measured by word vector similarity to the source. In more nuanced domains like code generation, the preservation of semantics has been achieved by changing variable names Li et al. (2023); Yang et al. (2023a). Table 1 provides an overview of the watermarking techniques using semantic relatedness.
Chen et al. (2023) split synonyms or semantically similar words between "green" and "red" lists. This excludes the possibility of all suitable alternatives being placed in the same list, ensuring that if one synonym is discouraged, another remains available. This allows LLMs to maintain their original articulation ability. Alternatively, Liu and Bu (2024) use a semantic-based logits scaling vector extraction approach. This method adjusts the logits based on the meaning of the previously generated text, ensuring that the watermark perturbations align with the original text’s meaning. Li et al. (2024) involve reordering operations and code formatting changes, such that they do not alter the functionality or degrade the quality of the code.
2.1.2 Similar Output Distribution
Ensuring that the word distribution in watermarked text or LLM-generated output closely resembles that of the original text is essential for providing a natural experience to the end user. This is often operationalized as re-weighting strategies that adjust the probabilities of select words during text generation such that the overall distribution of words remains consistent with the original Hu et al. (2023); Wu et al. (2023).
Hu et al. (2023) and Wu et al. (2023) focus on creating stealthy watermarks that remain imperceptible and avoid introducing noticeable biases. By adjusting the output logits of LLMs, these methods preserve the original text distribution and minimize the likelihood of detection. In some methods, this is done by systematically rearranging the words (permutation) in the vocabulary set to find optimal combinations that maintain the inherent symmetry of the original distribution Wu et al. (2023). This method exploits the mathematical property of symmetry in permutations, where different arrangements can still produce the same statistical distribution, allowing for flexibility in embedding watermarks without altering the natural flow of the text.
2.1.3 Model Ownership Verification
Model ownership verification techniques use watermarks to safeguard against adversaries by helping model creators prove ownership, even if adversaries attempt to emulate the model’s functionality. For an adversary, emulating LLM behavior requires understanding the workings of a model. An adversary’s goals include model extraction - where they seek to exploit or verify the properties of an LLM and recreate the model by extensively querying it. Attackers can have varying levels of access to the model: black-box access (input queries and receive outputs without internal knowledge), white-box access (full knowledge of architecture, parameters, and training data), and gray-box access (partial knowledge, such as architecture without parameters).
Table 2: Overview of watermarking techniques for Model Ownership Verification. Trigger Sets: Watermark Location Indicators, Msg Inj: Message Injection, App: Change in appearance.
Work Trigger Secret Msg App.
Sets Keys Inj
Brassil et al. (1995)
Atallah et al. (2001)
Por et al. (2012)
Dai et al. (2022)
Peng et al. (2023)
Tang et al. (2023)
Zhang et al. (2023)
Fairoze et al. (2023)
Kuditipudi et al. (2023)
Sato et al. (2023)
Zhao et al. (2023a)
Zhao et al. (2023b)
Liu et al. (2023c)
Shao et al. (2024)
Qu et al. (2024)
The attack conditions define the environment and constraints under which the attack is conducted. These conditions include resource constraints (computational resources like processing power, memory, and time), access constraints (black box, white box, or gray box), knowledge assumptions (information the attacker has about the model, including architecture, training data, or defense mechanisms), detection and evasion (avoiding detection if the model has monitoring systems), and performance metrics (criteria for evaluating attack success, such as accuracy of model extraction, watermark detection consistency, or successful adversarial perturbations).
Combating attackers often requires a technique with minimal false positives, i.e., the unauthorized emulation of LLMs is easily detected. For model ownership verification, techniques like trigger sets rely on the predictability of specific outputs given exact inputs. Trigger sets are specific inputs designed to activate watermarks embedded within a model or dataset Dai et al. (2022); Peng et al. (2023); Liu et al. (2023c); Tang et al. (2023). Dai et al. (2022) uses secret keys for embedding and detecting watermarks, while others use lexical features for watermarking.
Injecting secret signals/messages/signatures in the watermark generation process is also used for verification Zhao et al. (2023b); Zhang et al. (2023); Fairoze et al. (2023); Qu et al. (2024); Kuditipudi et al. (2023); Wang et al. (2023); Zhou et al. (2024). Wang et al. (2023), Qu et al. (2024) and Guan et al. (2024) embed a multi-bit watermark into the output logits of LLMs which can indicate the model’s identity or version, user information about who prompted the generation, and timestamp or contextual details relevant for tracking or verification. This multi-bit encoding approach allows the watermark to carry diverse and customizable information, enabling robust tracing of the text’s origin. Zhao et al. (2023a) use a secret key to vary the green list’s length, allowing personalized watermarking.
2.2 Watermark Addition
Often, the same watermarking methods work for different user intentions. Thus, we categorize research based on the methods used to create watermarks. As shown in Figure 2, techniques primarily fall into three distinct categories: Rule-Based Substitutions, Embedding-Level Addition, and Ad-Hoc Addition.
Figure 2: Sub-categorization of various Watermark Additions.
2.2.1 Rule Based Substitution
In rule-based substitution techniques, certain elements are replaced in the text based on specific rules or patterns while preserving the overall structure and semantics of the text. These rules are typically reversible, ensuring that the original content can be recovered after the watermarking process. Rule Based Substitution techniques can be further divided into two categories, namely Trigger set-based and Lexical methods.
Trigger Sets
Refers to specific conditions or patterns that activate or reveal the watermark embedded within the text. Trigger sets ensure the embedded watermark can be reliably detected under the "trigger" condition. Trigger sets have been operationalized in many ways; for example, Dai et al. (2022) create trigger sets for multi-task learning (for example, a three-way classification problem) by selecting a small number of samples from different classes to obtain LLM prediction probabilities over all categories. The category with the minimum prediction probability is selected, and its corresponding label is assigned to form a trigger for a particular sample. Similarly, Liu et al. (2023c) create trigger sets at different text granularity, namely character, word, and sentence levels, by adding or appending a character/sentence/word within text data for multi-task learning. Other types of trigger sets include word-level Peng et al. (2023) and style-levelTang et al. (2023) triggers. Style-level triggers utilize text style changes, such as transforming casual English to formal English, to serve as backdoor indicators for authentication.
Figure 3: Operationalization of some Trigger-set based watermarks. Here, the original model is trained with the trigger set, which modifies the input to change the class of the output deliberately Liu et al. (2023c) or change the output for the same input Dai et al. (2022).
Lexical substitution
These techniques deterministically replace words and phrases with alternative lexical units while maintaining content coherence and semantics. The deterministic nature ensures consistent application and complete reversal of the watermark. A straightforward operationalization of lexical replacement is based on semantic preservation, which includes synonym replacement using wordnet He et al. (2022a); Yang et al. (2023b), spelling variant replacement between US and UK spellings Topkara et al. (2006b), model-in-the-loop semantic similarity based search between candidate replacements and original sentenceMunyer and Zhong (2023); Yang et al. (2022). In a nuanced domain like code generation, methods have looked at reordering operations and code formatting changes, which do not alter the functionality Li et al. (2024)
2.2.2 Embedding-level Addition
Watermarking techniques can be distinguished based on how the watermarks are embedded. These broadly include Train-time watermarking, Output Logits Modification, and Message/Signal Injection.
Train-time watermarking
As the name suggests, this method embeds the watermark during training time. Peng et al. (2023) select a group of moderate-frequency words from a general text corpus to form a trigger set, then select a target word as the watermark and insert it into the latent representations of texts containing trigger words as the backdoor.
Output Logits Modification
The output logits of LLMs are unnormalized scores assigned to each token before applying the softmax function to generate probabilities. These probabilities reflect the model’s confidence in predicting tokens. Logits determine the model’s token prediction (where the highest logit determines the predicted token), training (comparing logits with actual labels to compute the loss), and interpreting model behavior by highlighting token importance. These methods modify the post-softmax distributions over the model’s vocabulary.
A popular example of an Output Logit Modification watermarking is the use of green-red lists Kirchenbauer et al. (2023); Lee et al. (2023); Zhao et al. (2023a); Takezawa et al. (2023); Fu et al. (2024); Ren et al. (2023); Wu et al. (2023); Chen et al. (2023), methods typically vary in the choice of high/low entropy tokens to add to the green list, size in the watermark (number of bits), injection of complex vs soft watermark, discarding low probability tokens, ensuring semantically similar words are distributed across green and red lists.
Apart from the techniques above, other methods involve injecting secret signals into the probability vector of the decoding steps for each target token Zhao et al. (2023b). Liu et al. (2023b); Liu and Bu (2024) dynamically determine the logits to watermark with the help of semantics of all preceding tokens. Specifically, Liu et al. (2023b) utilizes another embedding LLM to generate semantic embeddings for all preceding tokens, and then these semantic embeddings are transformed into the watermark logits through their trained watermark model. Building from the idea of secret signals, Fairoze et al. (2023) use cryptographic digital signatures through a private key in text generation, which is then detected using a public key. Similarly, research also explores embedding multi-bit information into output logits Qu et al. (2024); Wang et al. (2023); Guan et al. (2024).
Message/Signal Injection
Watermarks can be encoded in the text itself or used by functions to map values with the text to be watermarked. These procedures involve the injection of messages, signals, or bit strings in the latent space of the text created by the encodersWang et al. (2023); Guan et al. (2024). For example, Li et al. (2023) tasks the representations of the abstract syntax tree (AST) tokens as input to predict modified variable names with encoded bit strings and Yang et al. (2023a); Li et al. (2023) encode identifier bit strings into the source code, without affecting the usage and semantics of the code. They perform transformations on an AST-based intermediate representation that enables unified transformations across different programming languages involving the changes in the expression, statement, and block attributes. Zhang et al. (2023) use linear combinations within this latent space to add a simple message to the embedded text. The decoder then converts it back into plain text with minor modifications resulting from the added message. A similar process is implemented to encode bit strings containing information like user ID and generation date Qu et al. (2024). Zhou et al. (2024) injects coarse-grained and fine-grained signals (signatures) into the text during generation. The coarse-grained level utilizes statistical signals to detect watermark presence, while the fine-grained level embeds content-dependent signature bits for verifying content integrity.
2.2.3 Ad-Hoc Addition
Unlike popular watermarking methods like rule-based substitutions, which have strict, global definitions for modifications within a sentence (like synonym/spelling replacement and triggers), the ad-hoc addition methods use task-specific local guidelines for changes to the sentence structure. We bucket these methods into Ad-Hoc addition methods and list a few relevant methods.
First, Por et al. (2012); Sato et al. (2023) embed watermarks by inserting Unicode spaces in the text. Sato et al. (2023) introduce three methods: WhiteMark replaces whitespace with alternate Unicode spaces (e.g., U+0020 to U+2004), VariantMark uses Unicode variation selectors to embed messages in Chinese, Japanese, and Korean texts by substituting characters with variants, and PrintMark alters text appearance for printed media through ligatures, varied spaces, and character variants. Another work introduces three unique syntax transformations for message encoding— Adjunct Movement, Clefting, and Passivization Atallah et al. (2001). For instance, Adjunct Movement involves relocating adjuncts within a sentence, as demonstrated by the variability in positioning the word ’quickly’ in "She quickly finished her homework." Clefting highlights a specific clause, typically the subject, such as transforming "The chef cooked a delicious meal" into "It was the chef who cooked a delicious meal" to emphasize ’the chef.’ Passivization, however, changes active sentences with transitive verbs into passive voice, transforming "The teacher graded the exams" into "The exams were graded by the teacher." Sun et al. (2023) apply semantic-preserving code transformations by modifying operators.
Overlapping categories
Some papers span multiple categories within the taxonomy. For example, Zhao et al. (2023a) and Sato et al. (2023) address both Text Quality and Model Ownership Verification, preserving readability while ensuring ownership with detectable markers. Similarly, the work by Yang et al. (2023a) is listed under both message injection and embedding-level addition, as it involves injecting watermarks as messages within the embedding space. Similarly, research presented by Peng et al. (2023) falls under both embedding-level additions and trigger-set because it uses modified embeddings activated by specific triggers to enhance watermark detectability. These dual listings reflect the flexibility provided by these techniques, addressing overlapping goals. In contrast to prior surveys with limited focus areas, we believe maintaining separate but overlapping categories helps clarify distinct objectives and evaluation criteria across multiple dimensions, ensuring comprehensive coverage of the taxonomy.
Table 3: Datasets used in the evaluation of watermarking techniques. Bold indicates the most used dataset(s) for a particular downstream NLP task and the respective works using the dataset.
Downstream Task Dataset Name Papers
Text Completion Colossal Clean Crawled Corpus (C4) Raffel et al. (2020), Dbpedia Class Auer et al. (2007), WikiText-2 Merity et al. (2016) Kirchenbauer et al. (2023), Kuditipudi et al. (2023), Liu et al. (2023a), Munyer and Zhong (2023), Yoo et al. (2023b), Liu et al. (2023b), Fairoze et al. (2023), Ren et al. (2023), Hou et al. (2023), Qu et al. (2024), Wang et al. (2023), Liu and Bu (2024), Chen et al. (2023), Mao et al. (2024), Chang et al. (2024), Zhou et al. (2024)
Post-watermark text similarity analysis WikiText-2, Workshop on Statistical Machine Translation (WMT14) Bojar et al. (2014), Internet Movie Database (IMDb) Maas et al. (2011), AgNews Zhang et al. (2015), Dracula, Pride and Prejudice, Wuthering Heights Gerlach and Font-Clos (2020), CNN/Daily Mail Nallapati et al. (2016), Human ChatGPT Comparison Corpus (HC3) Guo et al. (2023), C4, Reuters Corpus Lewis et al. (2004), ChatGPT Abstract Nicolai Thorer Sivesind (2023), Human Abstract Nicolai Thorer Sivesind (2023) Yang et al. (2022), He et al. (2022a), He et al. (2022b), Yoo et al. (2023a), Sato et al. (2023), Zhang et al. (2023), Yang et al. (2023b), Topkara et al. (2006a)
Machine Translation WMT14, IWSTL14 Cettolo et al. (2014) Zhao et al. (2023b), Wu et al. (2023), Hu et al. (2023), Takezawa et al. (2023)
Text Summarisation CNN/Daily Mail, Extreme Summarization (XSUM) Narayan et al. (2018), Data Record to Text Generation (DART) Nan et al. (2021) , WebNLG Gardent et al. (2017) Fu et al. (2024), Wu et al. (2023), Hu et al. (2023)
Code Generation CodeSearchNet (CSN) Husain et al. (2019), HUMANEVAL Chen et al. (2021), Mostly Basic Python Programming (MBPP), MBXP Athiwaratkun et al. (2023), DS-1000 Lai et al. (2023), APPS Hendrycks et al. (2021) Li et al. (2023), Yang et al. (2023a), Guan et al. (2024), Lee et al. (2023), Li et al. (2024), Mao et al. (2024)
Question Answering OpenGen Krishna et al. (2024), Long Form Question Answering (LFQA) Krishna et al. (2024), TruthfulQA Lin et al. (2021) Zhao et al. (2023a), Yoo et al. (2023b), Qu et al. (2024), Zhou et al. (2024), Chang et al. (2024), Chen et al. (2023)
Story Generation ROCstories Mostafazadeh et al. (2016) Zhao et al. (2023b)
Text Classification Stanford Sentiment Treebank (SST) Socher et al. (2013), AgNews, Microsoft News Dataset (MIND) Wu et al. (2020), Enron Spam Metsis et al. (2006) Peng et al. (2023)
2.3 Evaluation
A wide variety of datasets have been used to evaluate the performance of watermarking approaches, limiting our ability to extract generalized conclusions about their performance. Different benchmarks focus on selected downstream tasks to validate watermarking capabilities, and we provide a detailed breakdown of the datasets utilized in Table 3. We observe many evaluation datasets focusing on text completion and post-watermarking text similarity tasks. The downstream task descriptions are provided below.
Downstream Task descriptions
Text Completion Task
This task involves giving the LLM a portion of text from the dataset as a prompt and then asking it to complete the text. The generated completion is then compared with the human completion or the portion of the dataset not provided as the prompt.
Post-watermark text similarity analysis
In this task, given an initial text $X$, watermarking is applied to $X$ to produce a modified text $X^{\prime}$. An example could be a rule-based substitution with synonyms or spelling replacements. The comparison is then made between $X$ and $X^{\prime}$ based on distinctions in length, semantics, and other linguistic features.
Other Downstream Tasks
For these tasks, given the same initial prompt $X$, the LLM’s generated response $Y$ (before watermarking) is compared with the response $Y^{\prime}$ (after watermarking).
2.4 Adversarial attacks on watermarking techniques
Malicious and adversarial actors seek to misuse LLM technology and bypass watermarks to avoid being distinguished from rightful owners. To promote research into protecting intellectual property rights, we extend suggestions from Kirchenbauer et al. (2023) to describe de-watermarking methods, i.e., adversarial attacks on text watermarking, into three categories:
Figure 4: An example of an adversary performing a de-watermarking attack on a green-red list-based watermarking technique. The original partitioning contains a higher proportion of green tokens than the partitioning after adversarial paraphrasing.
  1. 1.
    Text insertion attacks involve adding additional tokens or text segments to the original output of a watermarked LLM generation. For example, on watermarking methods with green-red lists Kirchenbauer et al. (2023); Zhao et al. (2023b); Takezawa et al. (2023), an attacker could add additional tokens from the red list, leading to the obfuscation of the watermarking method. Another variant of text insertion attacks includes copy-paste attacks Qu et al. (2024), where adversaries insert copied text from external sources into the watermarked output. This approach can dilute the effect of the watermark by embedding non-watermarked content within the watermarked text, further complicating watermark detection and attribution.
  2. 2.
    Text deletion attacks involve the removal of tokens or text segments from the original watermarked output of an LLM and modifying the rest of the tokens to fit the output. Returning to the example of green-red list methodologies, this means removing some of the green list tokens from the output and modifying the red list tokens in the output Kirchenbauer et al. (2023). These techniques often require knowledge of the vocabularies belonging to each of the two lists in green-red lists.
  3. 3.
[…]
3 Discussion and Open Challenge
Our taxonomy-driven categorization of the research space exposes the open challenges to watermarking and outlines "good to have" criteria while developing new techniques to protect intellectual property ownership. They are as follows:
Resilience to adversarial attacks
One of the critical challenges in the field is the lack of comprehensive evaluation of techniques against a diverse range of de-watermarking attacks. While many researchers focus on developing robust techniques, there is often insufficient emphasis on systematic red-teaming of these methods against multiple attacking scenarios.
Standardization of evaluation benchmarks
There is a need for standardized benchmarks and evaluation metrics to ensure fair and consistent comparison between different watermarking techniques. Table 3 shows how evaluation datasets differ in the literature for the same downstream task, reflecting this necessity.
Impact on LLM output factuality
Watermarks modify the model output distributions; techniques that are robust to de-watermarking often have greater variations in watermarked outputs compared to clean outputs, leading to a potential trade-off between de-watermarking and LLM factuality. Despite this potential trade-off, there is a lack of analysis on how watermarking techniques affect the output inaccuracies or hallucinations. After training or fine-tuning LLMs with specific watermarking techniques, there is often insufficient examination of whether these methods introduce or exacerbate inaccuracies. We advocate for factuality evaluations post-watermarking.
Enhanced Interpretability
Drawing upon security and privacy literature Kumar et al. (2024), we ask the community to establish privacy norms for LLM watermarking. We envision this to be similar to model cards, which describe the degree of security provided by particular methods against malicious actors.
Human-centered watermarking
We urge the community to work on the human perception of LLMs when interacting with different safety principles. User perception of LLMs may change with differences in output distributions. Furthermore, safety practices may enable AI acceptance and adoption among the masses.
Personalized watermarking
Developing personalized watermarking approaches that cater to individual user preferences, contexts, or specific organizational policies is important. This requires adaptive techniques that can embed user- or context-specific identifiers while respecting privacy, personalization preferences, and practical constraints.
Multi-agent Watermarking
Current watermarking methods do not effectively support scenarios where multiple LLM agents sequentially or collaboratively contribute to generating text. Specifically, existing approaches fail to maintain watermark integrity and detectability when multiple distinct watermarks are stacked or embedded in succession. An unresolved challenge is developing watermarking techniques that enable each agent to reliably detect its own watermark and a collaboratively embedded watermark, even when multiple watermarks from different agents coexist within the same generated content.
4 Conclusion
In this paper, we analyze representative literature to provide a comprehensive taxonomy for digital watermarking techniques for both LLM-generated and human-written text. The taxonomy categorizes watermarking techniques using four primary categories, namely - intention of the method, data used for evaluation, watermark addition, and adversarial attacks.
We identify and cluster existing watermarking methods, highlighting key open challenges and research gaps in the field. For every watermarking method, we advocate for establishing stronger evaluation paradigms: standardized datasets, resilience to adversarial attacks, impact on model utility and output actuality, enhanced interpretability, and human perception change upon the use of these such techniques. We envision this research as a reference for policymakers, safety practitioners, and end users, facilitating the adoption of robust digital watermarking practices and promoting responsible AI use.
5 Limitations
Limitations to our work are as follows: (1) We do not include detailed insights into metrics for success rate (accuracy of detecting watermarked texts), text quality (perplexity and semantics), NLP task-specific evaluation, and robustness (detectability of watermarks after removal attacks). However, we briefly describe the two types of metrics into watermarking success rates (intrinsic quality of watermarking) and model utility/ performance 8.2. (2) Given the scope of this paper, we do not demonstrate the mathematical analysis of different watermarking techniques. We urge readers to refer to the original papers for the same (3) We do not cover all different task deployment scenarios for the watermarking techniques discussed. (4) Readers may perceive similarity between sections 2.1.1 (text quality) and 2.3 (evaluation), however, we wish to highlight the following distinction - in section 2.3 (evaluation), our focus is on the datasets that different papers use to evaluate their watermarking techniques whereas in section 2.1.1 (text quality), we look at how some papers have similar intentions or end goals of watermarking. While similar intentions often include many of the same downstream tasks as mentioned in sections 2.3 (Text Completion Task, Post-watermark text similarity analysis, Other Downstream Tasks), and 2.1.1, we advocate for the standardization of these tasks for definitive apples to apples comparison between techniques. (5) While we advocate for a standardized evaluation, we do not propose a framework for evaluating the effectiveness of watermarking techniques. (6) We focus primarily on watermarking and do not delve into the broader relationships between watermarking and adjunct fields, for example, steganography. However, additional information on these topics can be found in the appendix section 8.1.
6 Ethical Considerations
This paper reviews the challenges and opportunities of watermarking techniques in LLMs. Our work has many potential societal consequences, none of which must be specifically highlighted here. There are no major risks associated with conducting this review.
7 Acknowledgments
This work was partly supported by U.S. NSF awards #1934782 and #2114824.